The width of the soccer field recommended for players under the age of 12 is
35 yd less than the length. The perimeter of the field is 330 yd. Find the dimensions of the field.

Answer by checkley77(12844)

W=L-35
2L+2W=330
2L+2(L-35)=330
2L+2L-70=330
4L=330+70
4L=400
L=400/4
L=100 YDS. IS THE LENGTH.
W=100-35=65 YDS. IS THE WIDTH.
PROOF:
2*100+2*65=330
200+130=330
330=330